# Diffusion Posterior Proximal Sampling for Image Restoration

### v1

@@ 3.1 Proximal Sampling at Each Step
1:  $\mathbf{x}_{T}=\sqrt{\bar{\alpha}_{T}}\mathcal{A}^{T}\mathbf{y}+\sqrt{1-\bar{\alpha}_{T}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
3:     compute $\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})$ via (8)
@@
6:        $D_{i}$ = $\|\mathcal{A}\big{(}\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}{\mathbf{z}_{t}^{i}}-C_{1}\mathbf{x}_{t}\big{)}-C_{2}\mathbf{y}\|_{2}^{2}$
@@
9:     $\mathbf{x}_{t-1}\leftarrow\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}_{t}$
@@
To tackle the challenges posed by random sampling, in this paper, we propose to extract multiple candidate samples from the predicted distribution, and select the most proximal one to our anticipated target. Our motivation comes from the following idea: considering $\mathbf{x}_{0}^{*}$ is an ideal but unknown solution for the image restoration problem, the sample taken from the predicted distribution $\mathbf{x}_{t-1}\sim p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ should be close to posterior $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}^{*})$, as it models the desired reverse process. The mathematical formulation of our selection process is given by:
$\mathbf{x}_{t-1}=$ $\operatorname*{arg\,min}_{\mathbf{x}_{t-1}^{i}}\|\mathbf{x}_{t-1}^{i}-\mathbf{x}_{t-1}^{*}\|_{2}^{2}$ (9)
where $\mathbf{x}_{t-1}^{i}\sim p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y}),~{}i\in[0,n-1]$, $\mathbf{x}_{0}^{*}$ denotes the unknown solution for Eq. (5), and $\mathbf{x}_{t-1}^{*}$ is our proposed deterministic solution via DDIM (Song et al., 2021a) (see Appendix B for detailed derivation):
$\mathbf{x}_{t-1}^{*}=$ $\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}^{*}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{*}}{\sqrt{1-\bar{\alpha}_{t}}}$ (10)
$=$ $C_{1}\mathbf{x}_{t}+C_{2}\mathbf{x}_{0}^{*}.$
where $C_{1}={\sqrt{1-\bar{\alpha}_{t-1}}}/{\sqrt{1-\bar{\alpha}_{t}}}$ and $C_{2}=\sqrt{\bar{\alpha}_{t-1}}-{\sqrt{\bar{\alpha}_{t}}\sqrt{1-\bar{\alpha}_{t-1}}}/{\sqrt{1-\bar{\alpha}_{t}}}$ are introduced for simplicity.
However, accessing $\mathbf{x}_{0}^{*}$ during inference is infeasible, rendering $\mathbf{x}_{t-1}^{*}$ also unknown. Now, one important contribution of this paper is to project Eq. (9) onto the measurement subspace. This is feasible as $\mathcal{A}\mathbf{x}_{t-1}^{*}$ can be approximated under the condition $\mathbf{y}\approx\mathcal{A}\mathbf{x}_{0}^{*}$.
$\mathcal{A}\mathbf{x}_{t-1}^{*}\approx C_{1}\mathcal{A}\mathbf{x}_{t}+C_{2}\mathbf{y}.$ (11)
$\mathbf{x}_{t-1}=\operatorname*{arg\,min}_{\mathbf{x}_{t-1}^{i}}\|\mathcal{A}(\mathbf{x}_{t-1}^{i}-\mathbf{x}_{t-1}^{*})\|_{2}^{2}$ (12)
Since $\mathbf{x}_{0}^{*}$ is expected to be more accurate than the denoised result $\hat{\mathbf{x}}_{0|t}$, the measurement $\mathbf{y}\approx\mathcal{A}\mathbf{x}_{0}^{*}$ can provide a stable and strong supervisory signal to correct the sampling result. Then by means of sample selection, we effectively control the result $\mathbf{x}_{t-1}$ within a more desirable region, even though we have no access to the $\mathbf{x}_{0}^{*}$.
As $\mathcal{A}$ is irreversible and complex in numerous scenarios, we propose to randomly sample $n$ candidates $\mathbf{x}_{t-1}^{i},i\in[0,n-1]$ from the predicted distribution $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$, and choose the one with the minimal deviation from our anticipated values. Finally, Eq. (9) becomes:
$\mathbf{z}_{t}=$ $\operatorname*{arg\,min}_{\mathbf{z}_{t}^{i}}\|\mathcal{A}\left(\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}_{t}^{i}-C_{1}\mathbf{x}_{t}\right)-C_{2}\mathbf{y}\|_{2}^{2}$ (13)
The process of sample selection is also denoted as noise selection, given by $\mathbf{x}_{t-1}^{i}=\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}_{t}^{i}$. We detail the full algorithm in Algorithm 2, and name our algorithm Diffusion Posterior Proximal Sampling (DPPS). The chosen $\mathbf{x}_{t-1}$ is referred to as the proximal sample due to the similarity to the proximal operator (Parikh & Boyd, 2014). An ablation study concerning the number of candidate samples $n$ is presented in Section 5.4.
Figure 3: Conceptual illustration of our DPPS. Our method extract multiple candidate samples from the predicted distribution and choose the one with the highest measurement consistency.
3.2 Aligned Initialization
When training, the diffusion model is trained with samples from $q(\mathbf{x}_{T}|\mathbf{x}_{0})$. However, diffusion-based restoration algorithm starts from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. Recent studies have pointed out that initializition have a significant impact on the generated results (Singh et al., 2022; Cao et al., 2023; Ma et al., 2023), and the discrepancy between the two distributions account for a discretization error (Benton et al., 2023). In this paper, inspired by (Everaert et al., 2023), we simply initialize the sample in the same way as during training, making the best use of the available measurement
$\mathbf{x}_{T}=\sqrt{\bar{\alpha}_{T}}\mathcal{A}^{T}\mathbf{y}+\sqrt{1-\bar{\alpha}_{T}}\bm{\epsilon}\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (14)
3.3 Discussion
@@
Relevance to Existing Methods.  DPS (Chung et al., 2023a) proposes to estimate the intractable likelihood under Laplacian approximation, and adopts random sampling from the predicted distribution. However, the randomly chosen sample may not be well-coordinated with the measurement information. Another study MCG (Chung et al., 2022) iteratively applies projection onto the measurement subspace after the denoising step to ensure data consistency. However, the imposed projections lead to accumulated error due to measurement noise.
@@
Computational Efficiency.  Selecting the proximal sample introduces some extra computational overhead. In practice, the diffusion model conducts back-propagation only once, aligning with the mainstream approaches (Song et al., 2022; Chung et al., 2022, 2023a; Song et al., 2023c). The evaluations for Eq. (13) are considerably cheaper than gradient calculation, making the additional computational costs manageable. We show the computational resources for different settings in Section C.2.
Theoretical Analysis.
Here, we provide some theoretical supports for our methodology. Despite our proposed proximal sampling being random, we can theoretically approach the desired point within an upper bound with mild assumptions. The proposition, detailed in Appendix B, proves that we can converge to our desired target $\mathbf{x}_{t-1}^{*}$, when the number of candidate samples is large enough. Furthermore, our method doesn’t require an exceptionally large number of candidates to achieve promising performance. This is because the variance of our method decreases when the proximal one is selected from any number of candidates.
@@
$f(\mathbf{z})=\|\mathcal{A}(\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}-C_{1}\mathbf{x}_{t})-C_{2}\mathbf{y}\|^{2}_{2}.$ (15)
@@
We provide a proof in Appendix B. In line with our theoretical analysis, empirical observations suggest that the proximal sample strategy demonstrates improved data consistency and reduced stochasticity, as evidenced by the experimental results.
Figure 4: Image restoration results with $\sigma_{y}=0.01$. Row 1: SR ($\times$4), Row 2: 80% inpainting, Row 3: Gaussian deblurring.
4 Related Work
To solve image restoration or image inverse problems with diffusion models, plenty of notable works (Dhariwal & Nichol, 2021; Kawar et al., 2022; Chung et al., 2023a; Song et al., 2023b) have been introduced. The first category of approaches (Saharia et al., 2021; Dhariwal & Nichol, 2021; Saharia et al., 2022) involves training conditional diffusion models or approximating the likelihood with synthetic image pairs as training data. However, these methods require specific training for each task and lack generalization across a wide range of inverse problems.
In contrast, the second category of approaches addresses inverse problems with unconditional diffusion models by guiding the reverse diffusion process (Song et al., 2022; Zhu et al., 2023). Several studies (Song et al., 2023b; Chung et al., 2023a; Song et al., 2023c) concentrate on estimating the time-dependent likelihood $p_{t}(\mathbf{y}|\mathbf{x}_{t})$ for posterior sampling. Among them, Song et al. (2023c) enhances the likelihood approximation with a Monte-Carlo estimate. Ma et al. (2023) tackles image super-resolution by selecting the best starting point from multiple choices. Recently, latent diffusion models have also seen advancements (Rombach et al., 2022; Vahdat et al., 2021) and have been widely adopted in various image inverse problem scenarios (Rout et al., 2023b; He et al., 2023; Song et al., 2023a; Chung et al., 2023b; Rout et al., 2023a).
5 Experiments
5.1 Setup
Dataset, Model.  To showcase the effectiveness of the proposed methods, we conducted experiments on two standard datasets, namely FFHQ 256$\times$256 (Karras et al., 2019) and ImageNet 256$\times$256 (Russakovsky et al., 2015). The evaluation encompasses the first 1k images in the validation set of each dataset. The diffusion models pre-trained on ImageNet and FFHQ are sourced from Dhariwal & Nichol (2021) and Chung et al. (2023a), respectively. For comprehensive comparisons, we include state-of-the-art diffusion-based image restoration solvers, including the Plug-and-play alternating direction method of multipliers (PnP-ADMM) (Chan et al., 2016), Score-SDE (Song et al., 2022), Denoising Diffusion Restoration Models (DDRM) (Kawar et al., 2022), Manifold Constrained Gradient (MCG) (Chung et al., 2022), DPS (Chung et al., 2023a), and DiffPIR (Zhu et al., 2023). To ensure fair comparisons, we utilized the same pre-trained diffusion models for all diffusion-based methods. Additional details, including hyper-parameters and metrics, are provided in Appendix D. All experiments were executed with a fixed random seed unless explicitly stated.
Table 1: Quantitative evaluation of image restoration tasks on FFHQ 256$\times$256-1k and ImageNet 256$\times$256-1k with $\sigma_{y}=0.01$. Bold: best, underline: second best.
Inpaint (random) Deblur (Gaussian) Deblur (motion) SR ($\times$ 4)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
FFHQ
PnP-ADMM (Chan et al., 2016) 27.32 0.743 0.349 63.19 25.97 0.744 0.260 94.50 25.86 0.724 0.278 59.06 26.94 0.792 0.292 90.11
Score-SDE (Song et al., 2021b) 21.91 0.610 0.371 58.19 18.37 0.256 0.629 169.68 15.79 0.246 0.635 176.73 24.10 0.705 0.363 75.50
MCG (Chung & Ye, 2022) 24.59 0.711 0.265 29.31 19.67 0.549 0.497 85.85 18.00 0.513 0.604 91.73 24.02 0.698 0.321 52.88
DDRM (Kawar et al., 2022) 21.85 0.643 0.378 78.40 26.38 0.743 0.298 62.72 - - - - 27.31 0.806 0.248 52.49
DPS (Chung et al., 2023a) 27.29 0.816 0.182 23.11 26.04 0.742 0.228 30.52 25.24 0.726 0.261 33.47 26.55 0.763 0.237 34.50
DiffPIR (Zhu et al., 2023) 27.96 0.820 0.210 30.38 25.38 0.700 0.276 32.00 22.74 0.654 0.331 83.42 24.74 0.710 0.273 33.03
Ours 27.47 0.834 0.168 18.98 26.40 0.752 0.220 29.29 26.88 0.783 0.214 28.95 26.93 0.778 0.203 26.30
ImageNet
PnP-ADMM (Chan et al., 2016) 25.14 0.685 0.405 66.54 21.97 0.584 0.419 63.15 21.86 0.676 0.408 61.46 23.95 0.710 0.346 71.41
Score-SDE(Song et al., 2021b) 16.25 0.241 0.653 102.56 21.31 0.594 0.467 82.54 13.56 0.418 0.661 89.62 17.69 0.316 0.624 96.67
MCG (Chung & Ye, 2022) 23.21 0.661 0.324 44.09 12.31 0.405 0.647 109.45 18.32 0.463 0.633 99.26 17.08 0.292 0.538 85.91
DDRM (Kawar et al., 2022) 19.34 0.460 0.555 147.00 23.67 0.635 0.401 66.99 - - - - 25.49 0.729 0.319 54.77
DPS (Chung et al., 2023a) 25.65 0.780 0.240 29.04 19.65 0.487 0.422 65.35 18.79 0.465 0.458 77.29 23.88 0.667 0.335 42.83
DiffPIR (Zhu et al., 2023) 25.85 0.797 0.235 33.16 22.03 0.569 0.395 54.71 19.86 0.534 0.433 79.23 24.78 0.712 0.302 39.25
Ours 24.97 0.758 0.217 24.90 22.70 0.609 0.364 51.21 21.65 0.592 0.375 51.35 24.79 0.712 0.276 33.75
Degradation Operator.  The degradation operators are defined as follows: (i) For inpainting, 80% of the pixels (all RGB channels) in the image are masked. (ii) For Gaussian blur, a blur kernel of size $61\times 61$ with a standard deviation of $3.0$ is employed. (iii) For motion blur, the kernel, following the procedure outlined in Chung et al. (2022), has a size of $61\times 61$ and an intensity value of $0.5$. (iv) For SR ×$\times~{}4$, the operator involves 4×4$\times$4 × bicubic down-sampling. Additive Gaussian noise with a variance of $\sigma_{y}=0.01$ is applied for all degradation processes.
5.2 Results
Main Results.  We present the statistic results of the general image restoration tasks on both FFHQ and ImageNet datasets, as detailed in Table 1. To the dataset with homogeneous scenarios, i.e., FFHQ, our proposed method demonstrates impressive results across all metrics, establishing its superiority over existing state-of-the-art methods. When dealing with more varied images within the ImageNet dataset, our approach exhibits substantial outperformance against all baseline methods in terms of FID and LPIPS, while maintaining comparable levels in PSNR.
Refer to captionGround TruthInputDPSOurs ($n$=2)Ours ($n$=10)Ours ($n$=100)
Figure 5: Visual results on SR ($\times$4) task to demonstrate the efficacy of our proposed method and to explore the impact of $n$. More results are presented in Appendix F.
The visual results for inpainting, SR, and Gaussian deblurring are shown in Figure 4, showcasing the evident superiority of our proposed method. While DDRM attains commendable distortion results with fewer Neural Function Evaluations (NFEs), it faces limitations in reliably restoration results for image inpainting tasks characterized by a very low rank of the measurement. DiffPIR achieves satisfactory results in various scenarios; however, its performance is tied to the analytical solution for the data consistency term and exhibits sensitivity to measurement noise. DPS differs from our approach in that it introduces random noise at each generation step, leading to over-smoothed and unstable restoration results, as depicted in Figure 4 (row 3). Conversely, our proposed method circumvents such drawbacks. The generative process is stabilized by directing the sample to a predefined target, yielding realistic and detailed restoration outcomes. It is noteworthy that the generated results from our method exhibit minimal variation with different seeds (refer to Appendix F for additional visual results), aligning with the intended design of identity preserving.
5.3 The Effect of Proximal Sampling
@@
Faster Convergence. To investigate the impact of our proximal sampling on the generated results, we conducted further experimental analyses using the first 150 images from the FFHQ validation set. In addition to the native DPS, we incorporated the results of DPS with DDIM (Song et al., 2021a) deterministic sampling, denoted as DPS_DDIM, which mitigates the effects of randomness. Results for (a) mean of $\|\mathbf{y}-\mathcal{A}\hat{\mathbf{x}}_{0|t}\|_{2}^{2}$, (b) average LPIPS of $\hat{\mathbf{x}}_{0|t}$, and (c) average PSNR of $\hat{\mathbf{x}}_{0|t}$ over timesteps are reported in LABEL:fig:Convergence_speed. It is evident that our method facilitates a more stable optimization process, yielding markedly superior results compared to DPS and DPS_DDIM within the same period. This observation also supports the claim that the measurement can provide reliable supervisory signals for the sampling result.

### v2

@@ 3.4. Aligned Initialization
1:  $\mathbf{x}_{T}=\sqrt{\bar{\alpha}_{T}}\mathcal{A}^{T}\mathbf{y}+\sqrt{1-\bar{\alpha}_{T}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
3:     compute $\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})$ via (8)
@@
6:        $D_{i}$ = $\|\mathcal{A}\big{(}\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}{\mathbf{z}_{t}^{i}}-C_{1}\mathbf{x}_{t}\big{)}-C_{2}\mathbf{y}\|_{2}^{2}$
@@
9:     $\mathbf{x}_{t-1}\leftarrow\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}_{t}$
@@
Refer to captionInputDDRMDPSDiffPIROursReference
Figure 4. Image restoration results with $\sigma_{y}=0.01$. Row 1: SR ($\times$4), Row 2: 80% inpainting, Row 3: Gaussian deblurring.
[…]
Image restoration results with $\sigma_{y}=0.01$. Row 1: SR ($\times$4), Row 2: 80% inpainting, Row 3: Gaussian deblurring.
3.5. Discussion
3.5.1. Relevance to Existing Methods.
@@ 3.5. Discussion
DPS (Chung et al., 2023a) proposes to estimate the intractable likelihood under Laplacian approximation, and adopts random sampling from the predicted distribution. However, the randomly chosen sample may not be well-coordinated with the measurement information. Another study MCG (Chung et al., 2022) iteratively applies projection onto the measurement subspace after the denoising step to ensure data consistency. However, the imposed projections lead to accumulated error due to measurement noise.
@@
3.5.2. Computational Efficiency.
Selecting the proximal sample introduces some extra computational overhead. In practice, the diffusion model conducts back-propagation only once, aligning with the mainstream approaches (Song et al., 2022; Chung et al., 2022, 2023a; Song et al., 2023c). The evaluations for Eq. (13) are considerably cheaper than gradient calculation, making the additional computational costs manageable. We show the computational resources for different settings in Section 5.4.
3.5.3. Theoretical Analysis.
Here, we provide some theoretical supports for our methodology. Despite our proposed proximal sampling being random, we can theoretically approach the desired point within an upper bound with mild assumptions. The proposition, detailed in supplementary material, proves that we can converge to our desired target $\mathbf{x}_{t-1}^{*}$, when the number of candidate samples is large enough. Furthermore, our method doesn’t require an exceptionally large number of candidates to achieve promising performance. This is because the variance of our method decreases when the proximal one is selected from any number of candidates.
@@
(16) $f(\mathbf{z})=\|\mathcal{A}(\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}-C_{1}\mathbf{x}_{t})-C_{2}\mathbf{y}\|^{2}_{2}.$
The variance for $f(\mathbf{z})$ is denoted by $\text{Var}\left(f(\mathbf{z})\right)$. We have
(17) $\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)>\text{Var}_{\text{MC}}\left(f(\mathbf{z})\right)>\text{Var}_{\text{Ours}}\left(f(\mathbf{z})\right).$
Here, $\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)$ is the variance of DPS (Chung et al., 2023a), $\text{Var}_{\text{MC}}\left(f(\mathbf{z})\right)$ is the variance of Monte Carlo sampling, and $\text{Var}_{\text{Ours}}\left(f(\mathbf{z})\right)$ is the variance of our proximal sampling method.
In line with our theoretical analysis, empirical observations suggest that the proximal sample strategy demonstrates improved data consistency and reduced stochasticity, as evidenced by the experimental results.
Table 1. Quantitative evaluation of image restoration tasks on FFHQ 256$\times$256-1k and ImageNet 256$\times$256-1k with $\sigma_{y}=0.01$. Bold: best, underline: second best.
Inpaint (random) Deblur (Gaussian) Deblur (motion) SR ($\times$ 4)
Method PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
FFHQ
PnP-ADMM (Chan et al., 2016) 27.32 0.349 63.19 25.97 0.260 94.50 25.86 0.278 59.06 26.94 0.292 90.11
Score-SDE (Song et al., 2021b) 21.91 0.371 58.19 18.37 0.629 169.68 15.79 0.635 176.73 24.10 0.363 75.50
MCG (Chung and Ye, 2022) 24.59 0.265 29.31 19.67 0.497 85.85 18.00 0.604 91.73 24.02 0.321 52.88
DDRM (Kawar et al., 2022) 21.85 0.378 78.40 26.38 0.298 62.72 - - - 27.31 0.248 52.49
DPS (Chung et al., 2023a) 27.29 0.182 23.11 26.04 0.228 30.52 25.24 0.261 33.47 26.55 0.237 34.50
LGD-MC (Song et al., 2023c) 26.11 0.257 35.26 24.08 0.288 34.54 21.04 0.370 49.19 26.12 0.247 33.68
DiffPIR (Zhu et al., 2023) 27.96 0.210 30.38 25.38 0.276 32.00 22.74 0.331 83.42 24.74 0.273 33.03
Ours 27.50 0.161 18.65 26.42 0.221 28.93 27.82 0.197 28.28 26.94 0.201 25.98
ImageNet
PnP-ADMM (Chan et al., 2016) 25.14 0.405 66.54 21.97 0.419 63.15 21.86 0.408 61.46 23.95 0.346 71.41
Score-SDE(Song et al., 2021b) 16.25 0.653 102.56 21.31 0.467 82.54 13.56 0.661 89.62 17.69 0.624 96.67
MCG (Chung and Ye, 2022) 23.21 0.324 44.09 12.31 0.647 109.45 18.32 0.633 99.26 17.08 0.538 85.91
DDRM (Kawar et al., 2022) 19.34 0.555 147.00 23.67 0.401 66.99 - - - 25.49 0.319 54.77
DPS (Chung et al., 2023a) 25.65 0.240 29.04 19.65 0.422 65.35 18.79 0.458 77.29 23.88 0.335 42.83
LGD-MC (Song et al., 2023c) 24.06 0.316 40.95 20.32 0.423 62.79 19.07 0.461 78.79 22.78 0.390 59.61
DiffPIR (Zhu et al., 2023) 25.85 0.235 33.16 22.03 0.395 54.71 19.86 0.433 79.23 24.78 0.302 39.25
Ours 24.97 0.217 24.90 22.70 0.364 51.21 21.65 0.375 51.35 24.44 0.267 30.70
4. Related Work
4.1. Diffusion Models for Image Restoration
To solve image restoration or image inverse problems with diffusion models, plenty of notable works (Dhariwal and Nichol, 2021; Kawar et al., 2022; Chung et al., 2023a; Song et al., 2023b) have been introduced. The first category of approaches (Saharia et al., 2021; Dhariwal and Nichol, 2021; Saharia et al., 2022) involves training conditional diffusion models or approximating the likelihood with synthetic image pairs as training data. However, these methods require specific training for each task and lack generalization across a wide range of inverse problems. In contrast, the second category of approaches addresses inverse problems with unconditional diffusion models by guiding the reverse diffusion process (Song et al., 2022; Zhu et al., 2023). Several studies (Song et al., 2023b; Chung et al., 2023a; Song et al., 2023c) concentrate on estimating the time-dependent likelihood $p_{t}(\mathbf{y}|\mathbf{x}_{t})$ for posterior sampling. Among them, Song et al. (2023c) enhances the likelihood approximation with a Monte-Carlo estimate. Ma et al. (2023) tackles image super-resolution by selecting the best starting point. (Jin et al., 2024; Jiang et al., 2023; Hou et al., 2024) effectively constrain the results of each generative step, adeptly tackling low-light image enhancement and shadow removal.
Recently, latent diffusion models have also seen advancements (Rombach et al., 2022; Vahdat et al., 2021) and have been widely adopted in various image inverse problem scenarios (Rout et al., 2023b; He et al., 2023; Song et al., 2023a; Chung et al., 2023b; Rout et al., 2023a). Furthermore, while several deep learning-based proximal optimization algorithms have been proposed (Wei et al., 2022; Lai et al., 2023), our method stands out as the first to make the diffusion sampling as a proximal optimization process.
4.2. The Exposure Bias in Diffusion Models
The exposure bias (Ning et al., 2023b; Li et al., 2023) in diffusion models is described as the misalignment between the training input $\mathbf{x}_{t}$ and the reverse process input $\mathbf{x}_{t}$, which is essentially a mismatch between the predicted noise $\boldsymbol{\epsilon}_{\mathbf{\theta}}$ and the actual noise $\boldsymbol{\epsilon}$. Ning et al. (2023b) mitigate the exposure bias problem by perturbing the training input, rendering the network more robust to inaccurate inputs during generation. Ning et al. (2023a) propose dynamic scaling to correct the magnitude error of $\boldsymbol{\epsilon}_{\mathbf{\theta}}$ and improve sample quality. Li et al. (2023) present a novel solution by adjusting the corruption level of the current samples. This paper provides a new way to effectively alleviate the exposure bias problem to some extent by reducing the uncertainty of random sampling.
5. Experiments
5.1. Setup
Dataset, Model.
To showcase the effectiveness of the proposed methods, we conducted experiments on two standard datasets, namely FFHQ 256$\times$256 (Karras et al., 2019) and ImageNet 256$\times$256 (Russakovsky et al., 2015). The evaluation encompasses the first 1k images in the validation set of each dataset. The diffusion models pre-trained on ImageNet and FFHQ are sourced from Dhariwal and Nichol (2021) and Chung et al. (2023a), respectively. For comprehensive comparisons, we include state-of-the-art diffusion-based image restoration solvers, including the Plug-and-play alternating direction method of multipliers (PnP-ADMM) (Chan et al., 2016), Score-SDE (Song et al., 2022), Denoising Diffusion Restoration Models (DDRM) (Kawar et al., 2022), Manifold Constrained Gradient (MCG) (Chung et al., 2022), DPS (Chung et al., 2023a), LGD-MC (Song et al., 2023c), and DiffPIR (Zhu et al., 2023). Sampling frequency parameter $N_{max}$ is set to 50. To ensure fair comparisons, we utilized the same pre-trained diffusion models for all diffusion-based methods. All experiments were executed with a fixed random seed.
Refer to captionReferenceInputDPSOurs ($n$=10)Ours ($n$=20)Ours ($n$=100)
Figure 5. Visual results on SR ($\times$4) task to demonstrate the efficacy of our proposed method and to explore the impact of $n$.
@@ 5.1. Setup
Visual results on SR ($\times$4) task to demonstrate the efficacy of our proposed method and to explore the impact of $n$.
Degradation Operator.
The degradation operators are defined as follows: (i) For inpainting, 80% of the pixels (all RGB channels) in the image are masked. (ii) For Gaussian blur, a blur kernel of size $61\times 61$ with a standard deviation of $3.0$ is employed. (iii) For motion blur, the kernel, following the procedure outlined in Chung et al. (2022), has a size of $61\times 61$ and an intensity value of $0.5$. (iv) For SR ×$\times~{}4$, the operator involves 4×4$\times$4 × bicubic down-sampling. Additive Gaussian noise with a variance of $\sigma_{y}=0.01$ is applied for all degradation.
Metrics.
The metrics employed for the comparison encompass: Peak Signal-to-Noise Ratio (PSNR) as distortion metrics; Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) distance, and Frechet Inception Distance (FID) (Heusel et al., 2017) as perceptual metrics.
5.2. Main Results
We present the statistic results of the general image restoration tasks on both FFHQ and ImageNet datasets, as detailed in Table 1. To the dataset with homogeneous scenarios, i.e., FFHQ, our proposed method demonstrates impressive results across all metrics, establishing its superiority over existing state-of-the-art methods. When dealing with more varied images within the ImageNet dataset, our approach exhibits substantial outperformance against all baseline methods in terms of FID and LPIPS, while maintaining comparable levels in PSNR.
The visual results for inpainting, SR, and Gaussian deblurring are shown in Figure 4, showcasing the evident superiority of our proposed method. While DDRM attains commendable distortion results with fewer Neural Function Evaluations (NFEs), it faces limitations in reliably restoration results for image inpainting tasks characterized by a very low rank of the measurement. DiffPIR achieves satisfactory results in various scenarios; however, its performance is tied to the analytical solution for the data consistency term and exhibits sensitivity to measurement noise. DPS differs from our approach in that it introduces random noise at each generation step, leading to over-smoothed and unstable restoration results, as depicted in Figure 4 (column 3). Conversely, our proposed method circumvents such drawbacks. The generative process is stabilized by directing the sample to a predefined target, yielding realistic and detailed restoration outcomes. It is noteworthy that the generated results by our method exhibit minimal variation with different seeds (refer to supplementary material), aligning with the intended design of identity preserving.
5.3. The Effect of Proximal Sampling
Figure 6. FID comparison between random sampling and proximal sampling with different values of $\lambda_{t}$.
@@ 5.3. The Effect of Proximal Sampling
FID comparison between random sampling and proximal sampling with different values of $\lambda_{t}$.
Table 2. Comparision of computational resources and performance with different $n$ on FFHQ SR×$\times$ 4.
Setting Memory / Growth Speed / Growth LPIPS / Gain
$n$=1 2599 MB /  0.0 % 56.36 s /  0.0 % 0.235 /  0.0 %
$n$=2 2603 MB /  0.2 % 56.64 s /  0.5 % 0.221 /  6.0 %
$n$=10 2621 MB / 0.8 % 56.82 s /  0.8 % 0.207 /  11.9 %
$n$=20 2643 MB /  1.7 % 57.25 s /  1.5 % 0.203 /  13.6 %
adaptive $n$ 2714 MB /  4.4 % 56.86 s /  0.9 % 0.201 /   14.5%
(a) MSE
(b) LPIPS
(c) PSNR
Figure 7. Convergence speed analysis. (a) mean error $\|\mathbf{y}-\mathcal{A}\hat{\mathbf{x}}_{0|t}\|^{2}_{2}$, (b) average LPIPS of $\hat{\mathbf{x}}_{0|t}$, and (c) average PSNR of $\hat{\mathbf{x}}_{0|t}$ with timesteps respectively. Our method achieves a faster optimization process and better restoration performance.
@@
Convergence speed analysis. (a) mean error $\|\mathbf{y}-\mathcal{A}\hat{\mathbf{x}}_{0|t}\|^{2}_{2}$, (b) average LPIPS of $\hat{\mathbf{x}}_{0|t}$, and (c) average PSNR of $\hat{\mathbf{x}}_{0|t}$ with timesteps respectively. Our method achieves a faster optimization process and better restoration performance.
5.3.1. Faster Convergence.
@@
To investigate the impact of our proximal sampling on the generated results, we conducted further experimental analyses using the first 150 images from the FFHQ validation set. In addition to the naive DPS, we incorporated the results of DPS with DDIM (Song et al., 2021a) deterministic sampling, denoted as DPS_DDIM, which mitigates the effects of randomness. Results for (a) mean of $\|\mathbf{y}-\mathcal{A}\hat{\mathbf{x}}_{0|t}\|_{2}^{2}$, (b) average LPIPS of $\hat{\mathbf{x}}_{0|t}$, and (c) average PSNR of $\hat{\mathbf{x}}_{0|t}$ over timesteps are reported in Figure 7(c). It is evident that our method facilitates a more stable optimization process, yielding markedly superior results compared to DPS and DPS_DDIM within the same period. This observation also supports the claim that the measurement can provide reliable supervisory signals for the sampling result.
5.3.2. More Robust to Hyper-parameter.